Raymond Company produces a variety of bicycles. One of its plants produces two bicycles: a mountain model and a racing model. At the beginning of the year, the following data were prepared for this plant:

In addition, the following information was provided so that overhead costs could be assigned to each product:

Raymond Company uses activity-based costing to calculate product costs.

Required:

A. Calculate the per unit product cost for a mountain bike.
B. Calculate the per unit product cost for a racing bike.
C. Assume Armstrong Company adds 35% to the cost of a mountain bike and 40% to the

cost of a racing bike obtain the selling prices. Calculate the selling prices for a mountain bike and a racing bike.

Mountain

Racing

Quantity

220,000

125,000

Selling Price

$1,200

$1,000

Unit Prime Cost

$ 400

$ 450

Activity Name

Activity Driver

Activity Cost

Mountain

Racing

Machining

Machine hours

$20,000,000

250,000

250,000

Engineering

Engineering hours

$ 8,000,000

75,000

50,000

Packing

Packing orders

$ 3,500,000

50,000

125,000
Answer #1

mountain

racing

activity name

rate

activity

cost

activity

cost

machining

40

2,50,000

1,00,00,000

2,50,000

1,00,00,000

engineering

64

75,000

48,00,000

50,000

32,00,000

packing

20

50,000

10,00,000

1,25,000

25,00,000

total

1,58,00,000

1,57,00,000

Machining rate = 20,000,000 / 500,000 = 40 per machine hour

Engineering rate = 8,000,000 / 125,000 = 64 per engineering hour

Packing rate = 3,500,000 / 175,000 = 20 per packing order

particular

mountain

racing

unit prime cost

400

450

unit overhead cost

71.81

125.6

unit product cost

471.81

575.6

selling price

636.94

805.8